# Transitioning from perturbative to non-perturbative splittings within QCD jets

### v2

@@ IV.2 Jet selection based on RSD
Since we are interested in the charge ratio’s sensitivity to hadronization features and selecting jets based on jet substructure variables, we need to check the resilience of our findings against pQCD effects, which can be done by comparing the charge ratios obtained with different parton shower descriptions provided by the two generators. Fig. 10 shows Pythia’s charge ratios for $N_{RSD}/N_{SD}>0.5$ jets and Fig. 11 the $r_{c}$ from Herwig, with all other parton shower models available in these Monte Carlo. Therefore, Fig. 10 has the $r_{c}$ from Pythia simulations using the DIRE (full markers) and VINCIA (open markers) showers and Fig. 10 the $r_{c}$ from Herwig simulations with the default shower. The panels below have the ratios calculated with respect to the simple and dipole baselines we have been using for Pythia and Herwig, respectively. The plateau nature of the Pythia $r_{c}$ is shown to be independent of the parton shower model employed, with variations limited to less than 10% relative to the simple shower baseline. In particular, the variations of the $r_{c}$ for jets with leading kaons (red) and leading protons (blue) are smaller than 2%. For the Herwig charge ratios, we observe that there is only a slight dependence on parton shower models of the early-time $r_{c}$, where jets with leading charged pions have a slightly less random charge configuration, jets with leading kaons have a stronger correlation and jets with leading protons remain basically the same. We can therefore conclude these charge ratio results to be robust against changes in the perturbative description, isolating the $r_{c}$ sensitivity to npQCD effects.
Figure 11: Charge ratio with respect to the LCP formation time for jets with late-RSD and with leading charged pions (black), kaons (red) and protons (blue); results obtained with the default parton shower model from Herwig 7; bottom panel shows the ratio with respect to the $r_{c}$ from Herwig’s dipole model.
V V - Conclusions
@@ VI Acknowledgments
We thank Isaac Mooney, Laura Havener, Yang-Ting Chien, Youqi Song, Krishna Rajagopal and Jose Guilherme Milhano for their valuable discussions and comments on the manuscript. This work was supported by the European Council (ERC) project ERC-2018-ADG-835105 YoctoLHC; by OE Portugal, Fundação para a Ciência e a Tecnologia (FCT), I.P., project CERN/FISPAR/0032/2021. LA and NOM acknowledges the financial support by FCT under contract 2021.03209.CEECIND and under PhD grant PRT/BD/154611/2022, respectively. RKE would also like to acknowledge funding by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grant number DE-SC0024660.

### v3

@@ III.2 Jet selection based on RSD
Since we are interested in the charge ratio’s sensitivity to hadronization features within varying jet selections, we check the resilience of our findings against primarily changes in pQCD aspects of the models. In Appendix A, we show the charge correlation ratios for a fixed parton shower description, isolating the effects due to Lund string and cluster fragmentation models for hadronization, reproducing the results from Fig. 9 with minor differences. In addition, Figs. 10 and 11 quantify the impact of varying parton shower models on $r_{c}$. Fig. 10 shows Pythia’s charge ratios for $N_{RSD}/N_{SD}>0.5$, and similarly Fig. 11 the $r_{c}$ from Herwig, with all other parton shower options available in the event generation. We show the Pythia simulations using the DIRE (full markers) and VINCIA (open markers) showers in Fig. 10 and Fig. 11 the $r_{c}$ from Herwig simulations with the default shower. The bottom panels in each figure show the ratios with respect to the simple and dipole baselines we have been using for Pythia and Herwig, respectively. The plateau nature of Pythia’s $r_{c}$ is shown to be relatively independent of the parton shower model employed, with minimal variations limited to less than 10% relative to the simple shower baseline. In particular, the variations of the $r_{c}$ for jets with leading kaons (red) and leading protons (blue) are smaller than 2%. For the Herwig charge ratios, we observe that there is only a slight dependence on parton shower models of the early-time $r_{c}$, where jets with leading charged pions have a slightly less random charge configuration. Similarly, we find jets with leading kaons have a stronger correlation whereas jets with leading protons remain basically the same. We can therefore conclude these charge ratio results to be robust against changes in the perturbative description, isolating the $r_{c}$ sensitivity to npQCD effects.
Figure 11: Charge ratio with respect to the LCP formation time for jets with late-RSD and with leading charged pions (black), kaons (red) and protons (blue); results obtained with the default parton shower model from Herwig 7.2; bottom panel shows the ratio with respect to the $r_{c}$ from Herwig’s dipole model.
IV V - Conclusions
@@ V Acknowledgments
We thank Isaac Mooney, Laura Havener, Yang-Ting Chien, Youqi Song, Krishna Rajagopal and Jose Guilherme Milhano for their valuable discussions and comments on the manuscript and Andrzej Siodmok and Pratixan Sarmah for their help on accessing state-of-the-art Monte Carlo tools. This work was supported by the European Council (ERC) project ERC-2018-ADG-835105 YoctoLHC; by OE Portugal, Fundação para a Ciência e a Tecnologia (FCT), I.P., project CERN/FISPAR/0032/2021. LA and NOM acknowledges the financial support by FCT under contract 2021.03209.CEECIND and under PhD grant PRT/BD/154611/2022, respectively. RKE would also like to acknowledge funding by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grant number DE-SC0024660.
VI Appendix A: Further studies on the impact of shower and hadronization models
Since both Monte Carlo event generators provide more than one parton shower description for the pQCD dynamics, we compared the different options between Pythia and Herwig. In the top panel of Fig. 12, it is represented the transverse momentum ($p_{T}$) spectra of final-state particles Pythia using Lund string hadronization. The spectra are shown for three different parton showers: simple (full red circles), VINCIA (full blue circles), and DIRE (full black circles). Corresponding data from Herwig (cluster hadronization model) using both default (open blue circles) and dipole (open red circles) showers are also included. The middle and bottom panels display the ratios of Pythia’s spectra to those from Herwig’s simple and dipole showers, respectively. We can see that the simple shower from Herwig most closely aligns with Pythia’s output, especially in the dominant low-$p_{T}$ region.
Figure 12: Top panel: particle transverse momentum spectra ($p_{T}>200$ MeV/c) for Pythia’s simple shower (red, full), VINCIA (blue, full) and DIRE (black, full) and Herwig’s default shower (blue, open) and dipole (red, open); Middle panel: ratios of Pythia spectra with respect to Herwig default; Bottom panel: ratios of Pythia spectra with respect to Herwig dipole.
The rapidity ($y$) distributions were also assessed and are presented in Fig. 13 using the same panel arrangement and marker scheme as previously mentioned. This analysis reinforces the superior alignment of the dipole shower from Herwig with Pythia’s rapidity profiles, thus substantiating our decision to compare DIS simulations using Pythia’s simple shower and Herwig’s dipole shower.
Figure 13: Top panel: particle rapidity for Pythia’s simple shower (red, full), VINCIA (blue, full) and DIRE (black, full) and Herwig’s default shower (blue, open) and dipole (red, open); Middle panel: ratios of Pythia distributions with respect to Herwig default; Bottom panel: ratios of Pythia distributions with respect to Herwig dipole.
Furthermore, we isolated the influence of hadronization models by maintaining a consistent parton shower prescription (angular ordered from Herwig) and varying only the hadronization approach. Fig. 14 shows the $p_{T}$ spectra for final-state particles using Pythia’s Lund string model (full circle markers) and Herwig’s cluster model (open circle markers). This interface with Pythia is done via TheP8I the , available in the newest Herwig 7.3 version Bewick et al. (2024). The rapidity distributions, shown in Fig. 15, adhere to the same marker and panel configuration.
Figure 14: Top panel: particle transverse momentum spectra ($p_{T}>200$ MeV/c) for Pythia’s Lund string model (red, full) and Herwig’s cluster fragmentation model (red, open), both applied to Herwig’s dipole shower; Bottom panel: ratios of the Lund spectra with respect to the cluster ones.
Figure 15: Top panel: particle rapidity for Pythia’s Lund string model (red, full) and Herwig’s cluster fragmentation model (red, open), both applied to Herwig’s dipole shower; Bottom panel: ratios of the Lund spectra with respect to the cluster ones.
Given the small differences on the particle and jet yield differences with respect to the main text results, we aimed to further test the resilience of the qualitative discrepancies between the charge ratios from Pythia and Herwig for late RSD jets (see Fig. 9) against the perturbative description of the parton shower. As such, Fig. 16 shows the equivalent charge ratios using the Herwig angular ordered parton shower description, but with the full circle markers representing the Lund string model as hadronization mechanism, and the open circle markers the cluster model.
Figure 16: Charge ratios with respect to the LCP formation time for jets with leading charged pions (black), kaons (red) and protons (blue) computed using Herwig 7.3, where full circles correspond to the implementation of the Lund string model and the open circles to the cluster fragmentation model; top panel shows early RSD jets and bottom panel shows late RSD jets.
This assessment continues to highlight a significant divergence in charge ratios for early-time LCP and late RSD jet selections, despite mild numerical differences. Owing to a generally more consistent agreement between the event profiles generated by both models compared to the initial Herwig versus Pythia comparison, we expanded this analysis to include jets with $p_{T,jet}>5$ GeV/c, as suggested in Chien et al. (2022). The outcomes for the lower momentum jets are shown in Fig. 17, focusing on late RSD jets, which better align with the findings presented in the bottom panel of Fig. 9.
Figure 17: Charge ratios with respect to the LCP formation time for late RSD, $p_{T,jet}>5$ GeV/c jets with leading charged pions (black), kaons (red) and protons (blue) computed using Herwig 7.3, where full circles correspond to the implementation of the Lund string model and the open circles to the cluster fragmentation model.